OpenStudy (anonymous):

One slice of bread is 80 calorie. Approximately how many calories are in 2 ½ slices of bread? A. 140 calories B. 200 calories C. 220 calories D. 240 calories E. 260 calories

OpenStudy (anonymous):

2 1/2 times 80 = 5/2 times 80 = 200

OpenStudy (anonymous):

2 1/2 or 5/2 times 80/1 is 400 over 2 which equals 200 calories.